A light ray traveling through medium 1, index of refraction n1 = 1.80, reaches the interface between medium 1 and medium 2, index of refraction n2 = 1.24. At what minimum angle with respect to the normal must the ray be incident on the interface in order to be totally internally reflected? In format 10.0


1
Expert's answer
2021-05-05T10:54:34-0400

"\frac{\sin\alpha}{\sin\beta}=\frac{n_2}{n_1},"

"\beta=90°,~\sin\beta=1,"

"\alpha=\arcsin \frac{n_2}{n_1}=\arcsin\frac {1.24}{1.80}=43.5°."